# OpenDriveVLA: Towards End-to-end Autonomous Driving with Large Vision Language Action Model

### v1

@@ 4.1 Training Data
Table 2: Open-Loop planning performance comparison of different driving models, including both autoregressive methods and non-autoregressive methods. OpenDriveVLA shows powerful planning ability and achieves best-in-class results among open-source models, even with the 0.5B version. We refer to the result summary from [35, 29, 23, 15].
@@
Motion Forecasting and Trajectory Prediction. We formulate both agent motion forecasting and ego trajectory planning in the ego vehicle system, where the model directly predicts future displacements within each entity’s local coordinate frame relative to the ego vehicle for planning and relative to each agent for forecasting. This formulation captures motion dynamics in a spatially consistent manner across all entities. Following [29], the ego vehicle state is encoded as textual input to ensure ego awareness throughout the training process. Both tasks predict 3-second future trajectories, sampled at 0.5-second intervals, resulting in 6 waypoints per trajectory.
4.2 Evaluations
@@
We assess OpenDriveVLA on open-loop planning task on nuScenes benchmark, where the model is evaluated following both ST-P3 [13] and UniAD [14] settings. The evaluation metrics include L2 displacement errors at 1, 2, and 3 seconds, along with the average collision rate over the prediction horizon. To evaluate the scene understanding ability of OpenDriveVLA, we assess its performance on three driving VQA datasets directly after the driving instruction tuning stage, i.e., [46], nuScenesQA [32], and nuX [8]. The VQA evaluation adopts standard MLG metrics, including BLEU, METEOR, CIDEr, BERT-Score, etc.
4.3 Implementation Details
@@
The 3D visual perception module in OpenDriveVLA follows the vision-centric design from [14], using a ResNet-101 backbone for 2D feature extraction. The perception backbone is pre-trained via multi-task learning on 3D object detection, object tracking, and map segmentation. The resulting BEV feature map has a spatial resolution of $200\times 200$. To construct a unified scene representation, the global SceneSampler applies 2D adaptive pooling to each camera view, subsequently concatenating the pooled multi-view features into a global scene token. Agent and map tokens are extracted from the final layer of their respective QueryTransformer modules. Each token type is then mapped into the language space using a separate two-layer MLP with GeLU activation. We adopt Qwen 2.5-Instruct [45] as the pre-trained LLM, which undergoes full parameter tuning during training. The model is trained on 4 NVIDIA H100 GPUs with a batch size of 1, completed in approximately two days. During inference, we set the decoding temperature to 0 to ensure deterministic trajectory generation. We freeze the 2D backbone during stage 3. Detailed training configurations are summarized in Table 1.
Models CIDER BLEU-4 METEOR ROUGE-L
Hint-UniAD [8] 21.7 4.2 12.7 27.0
Hint-VAD [8] 22.4 4.2 13.2 27.6
GPT-4o [44] 19.0 4.0 10.3 24.9
Gemini 1.5 [36] 17.6 3.4 9.3 23.4
Vote2Cap-DETR [6] 15.3 2.6 10.9 24.2
TOD3Cap [20] 14.5 2.5 10.5 23.5
@@
Table 3: Performance comparison on the Nu-X dataset [8]. OpenDriveVLA outperforms prior works in most metrics.
Method nu-Caption nuScenes-QA
BL-1 BL-2 BL-3 BL-4 BERT-S Ext Cnt Obj Sts Cmp H0 H1 Acc
Mini-GPT4 [55] 15.0 6.8 3.7 2.6 84.4 - - - - - - - -
Instruct-BLIP [7] 18.7 13.4 7.4 5.2 85.9 - - - - - - - -
LLaMA-AdapV2 [9] 30.2 17.3 10.4 7.5 86.5 19.3 2.7 7.6 10.8 1.6 15.1 4.8 9.6
LLaVA1.5 [27] 20.0 12.1 8.6 5.4 85.0 45.8 7.7 7.8 9.0 52.1 25.7 41.5 26.2
LiDAR-LLM [47] 41.0 30.0 23.4 19.3 91.3 74.5 15.0 37.8 45.9 57.8 - - 48.6
BEVDet+BUTD [32] - - - - - 83.7 20.9 48.8 52.0 67.7 - - 57.0
Table 4: Comprehensive performance comparison across nu-Caption [46] and nuScenes-QA [32]. Here, BL-1/2/3/4 refer to BLEU scores at different n-gram levels, BERT-S refers to BERT-Score, and metrics under nuScenes-QA correspond to different question categories: Ext (Existence), Cnt (Counting), Obj (Object), Sts (Status), and Cmp (Comparison), with Acc representing overall accuracy.
4.4 Main Results
Open Loop Trajectory Planning. We evaluate OpenDriveVLA on the open-loop trajectory planning task using both ST-P3 and UniAD metrics, ensuring comprehensive performance assessment across spatial accuracy and collision avoidance. As shown in Table 2, OpenDriveVLA achieves state-of-the-art performance across both settings. Specifically, both 3B and 7B version OpenDriveVLA achieve an average L2 error of 0.33m under ST-P3 metrics, outperforming prior autoregressive language models [29, 38]. On the UniAD metrics, OpenDriveVLA-7B also achieves great performance with an average L2 error of 0.66m. These results validate the effectiveness of OpenDriveVLA. Notably, OpenDriveVLA-0.5B, despite having significantly fewer parameters, also achieves competitive performance and outperforms prior larger models by a substantial margin. This highlights the efficiency of OpenDriveVLA, enabling strong spatial and semantic reasoning even with reduced model size, making it an effective and scalable solution for language-guided planning.
Driving Question Answering. We evaluate OpenDriveVLA on the driving VQA task across three nuScenes-based datasets, as shown in Table 4 and Table 3. OpenDriveVLA reaches best-in-class performance across all three datasets, consistently outperforming previous language-enhanced driving models and general-purpose multimodal baselines among most metrics. On nuCaption dataset, OpenDriveVLA achieves the best captioning performance among all evaluated models, outperforming both general multimodal LLMs LLaVA1.5 [27] and Mini-GPT4 [55], as well as autonomous driving-specific models such as LiDAR-LLM [47]. For nuScenesQA dataset, OpenDriveVLA also achieves strong performance. Compared to models that directly fuse BEV features with language models such as BEVDet+BUTD [32], OpenDriveVLA demonstrates clear advantages in object and status-related questions, which highlights the benefit of its spatially grounded visual-language alignment. Notably, OpenDriveVLA-0.5B outperforms even the larger 7B on the Nu-X dataset, which shows its powerful scene-understanding ability even with lightweight LLMs.
4.5 Ablation Study
Our ablation study examines the influence of input modalities on OpenDriveVLA’s trajectory planning. The results in Table 5 show that visual inputs enhance the action-making process, while textual high-level commands and historical state information further refine trajectory generation, indicating the contribution of semantic intent and temporal context. Besides, we find that ego-state information plays a significant role in trajectory prediction on nuScenes open-planning benchmark, aligning with prior findings [25]. Due to the imbalanced distribution of the dataset, where many scenarios involve maintaining the current state, the model tends to over-rely on ego-state history, leading to a bias toward conservative predictions. Hence, to further evaluate the generalization and instructed action-following capability of OpenDriveVLA, we conduct qualitative studies to examine its ability to execute diverse actions in response to different human commands.
Visual Ego Hist Cmd Avg. L2 (m) ↓ Avg. Collision (%) ↓
@@ 4.5 Ablation Study
Table 5: Ablation study on the effect with different input combinations on OpenDriveVLA-0.5B model. The table shows the results under UniAD and ST-P3 metrics.
Figure 3: Visualization of OpenDriveVLA-7B planning actions under different driving instructions and the visual question answering results. The agent motion prediction results are visualized after the agent-env-ego interaction stage. QA showcases are from results in Table 3, 4.
Figure 4: Qualitative results comparison of open-loop planning. The top figure shows the planning results of UniAD [14], while the bottom figure presents those of OpenDriveVLA-7B (Ours). The agent motion prediction results are visualized after the agent-env-ego interaction stage.
4.6 Qualitative Results
@@
Driving Command Following. Figure 3 illustrates trajectory planning at an intersection under two different driver instructions: keep forward and turn right, with the right turn as the ground truth. The results demonstrate OpenDriveVLA’s ability to accurately interpret driver commands and generate corresponding trajectory actions. The predicted trajectories adhere to the given command while remaining environment-aware and contextually reasonable, demonstrating strong command-following ability and generalization capability of our model in complex driving scenes.
Comparison with Prior Methods. Figure 4 compares the open-loop planning results of OpenDriveVLA-7B with UniAD [14]. The qualitative results demonstrate that OpenDriveVLA generates more stable and adaptable trajectories, as UniAD tends to overreact to parked vehicles on the right in narrow-road scenarios. Compared to UniAD, OpenDriveVLA effectively maintains trajectory smoothness and environmental awareness, demonstrating its improved capability for handling complex driving scenarios.
4.7 Discussions and Limitation
While OpenDriveVLA performs well across benchmarks, some limitations remain. It lacks explicit CoT reasoning, relying instead on implicit reasoning from driving instruction tuning, which may weaken inference capabilities in complex scenarios. Besides, despite limited token inputs and outputs, its autoregressive nature hampers real-time deployment in high-speed driving, requiring further optimization. Additionally, its evaluation in an open-loop setting does not account for robustness in interactive traffic environments.
5 Conclusion

### v2

@@ III-C Stage 2 - Driving Instruction Tuning
TABLE I: Open-Loop planning performance comparison of different driving models, including both autoregressive methods and non-autoregressive methods. OpenDriveVLA shows powerful planning ability and achieves best-in-class results among open-source models, even with the 0.5B version. We refer to the result summary from [song2025insightdriveinsightscenerepresentation, mao2024gptdriver, GPVL, RDA-Driver].
III-E Stage 3 - End-to-end Trajectory Planning Tuning
In this stage, OpenDriveVLA predicts ego trajectories as discrete waypoint sequences within a short horizon, denoted as $\mathcal{W}_{ego}=\{w_{1},w_{2},\dots,w_{T}\}$. Each waypoint $w_{t}$ represents the 2D coordinates $(x_{t},y_{t})$ of the ego vehicle at time step $t$. The waypoints are tokenized into a sequence of discrete textual tokens for autoregressive generation in the LLM: $\mathcal{T}{traj}=\mathrm{Tokenizer}(\mathcal{W}_{ego})$. The generation process is then cast as a causal sequence prediction task, where each token is predicted in a causal manner, conditioned on the visual perception tokens $\textbf{V}_{env}$, the ego state $\textbf{S}_{ego}$, and the driving command $\textbf{X}_{dri}$.
$\hat{\mathcal{T}}_{traj}=\textit{argmax}_{\textbf{T}_{traj}}\prod_{t=1}^{T}p\left(w_{t}\mid w_{1:t-1},\textbf{V}_{env},\textbf{S}_{ego},\textbf{X}_{dri}\right)$
The entire pipeline, including the 3D visual encoder, cross-modality projectors, and LLM, is jointly optimized end-to-end during training, with the 2D encoder kept frozen. At inference, the model autoregressively generates the tokenized trajectory $\hat{\mathcal{T}}_{traj}$, which is then decoded back into numerical waypoints:
$\hat{\mathcal{W}}_{ego}=\mathrm{Decoder}(\hat{\mathcal{T}}_{traj})$ (6)
IV Experiments
IV-A Training Datasets
We curate the training data of OpenDriveVLA based on its distinct training phases, drawing from: TOD3Cap [TOD3Cap], nuCaption [nucaption], nuScenesQA [nuscenesqa], nuX [hintad-nux], and GPT-Driver [mao2024gptdriver]. We conduct experiments on nuScenes [nuScenes], following standard data split into training and validation sets. OpenDriveVLA is trained using the training set paired with corresponding QA captions, while the validation set is exclusively used for performance evaluation to ensure fair comparisons with prior works. The details of training data can be found in supplementary materials.
Hierarchical Vision-Language Alignment. For agent-level caption, we post-process data from [TOD3Cap], which provides the 2D visual description of individual objects. To further enhance spatial grounding, each object caption is augmented with its corresponding BEV coordinates, enabling the model to associate object attributes with precise spatial locations. For scene tokens, we process multi-view scene descriptions from [nucaption], merging them into unified summaries that describe the driving environment across all camera views. For map tokens, structured language descriptions are derived from ground-truth annotations, translating map elements such as lane dividers, crosswalks, and road boundaries into descriptive text.
Driving Instruction Tuning. We adopt multiple instruction-oriented datasets derived from nuScenes to inject driving-specific knowledge into OpenDriveVLA. We unify several datasets into a standardized instruction-based QA format, including driving-related question-answer pairs collected from nuCaption [nucaption], nuScenesQA [nuscenesqa], and nuX [hintad-nux] dataset. Each QA pair is conditioned on structured environmental visual tokens and the ego vehicle state, ensuring consistency across different data sources. This multimodal instruction tuning process allows OpenDriveVLA to effectively ground language understanding into both environmental perception and scene understanding, bridging perception, reasoning, and action within the language space.
@@ IV-A Training Datasets
Motion Forecasting and Trajectory Prediction. We formulate both agent motion forecasting and ego trajectory planning in the ego system, where the model directly predicts future displacements within each entity’s local coordinate frame relative to the ego vehicle for planning and relative to each agent for forecasting. This formulation captures motion dynamics in a spatially consistent manner across all entities. Following [mao2024gptdriver], the ego vehicle state is encoded as textual input to ensure ego awareness throughout the training process. Both tasks predict 3-second future trajectories, sampled at 0.5-second intervals, resulting in 6 waypoints per trajectory.
Method nu-Caption nuScenes-QA
BL-1 BL-2 BL-3 BL-4 BERT-S Ext Cnt Obj Sts Cmp H0 H1 Acc
Mini-GPT4 [zhu2024minigpt4] 15.0 6.8 3.7 2.6 84.4 - - - - - - - -
Instruct-BLIP [dai2023instructblip] 18.7 13.4 7.4 5.2 85.9 - - - - - - - -
LLaMA-AdapV2 [gao2023llama-adapter-v2] 30.2 17.3 10.4 7.5 86.5 19.3 2.7 7.6 10.8 1.6 15.1 4.8 9.6
LLaVA1.5 [llava1_5] 20.0 12.1 8.6 5.4 85.0 45.8 7.7 7.8 9.0 52.1 25.7 41.5 26.2
LiDAR-LLM [nucaption] 41.0 30.0 23.4 19.3 91.3 74.5 15.0 37.8 45.9 57.8 - - 48.6
BEVDet+BUTD [nuscenesqa] - - - - - 83.7 20.9 48.8 52.0 67.7 - - 57.0
TABLE II: Performance on nu-Caption [nucaption] and nuScenes-QA [nuscenesqa]. BL-1/2/3/4: BLEU scores. QA metrics report accuracy on five question types: Existence, Counting, Object, Status, and Comparison.
IV-B Evaluations
@@
We evaluate OpenDriveVLA on the open-loop planning task of nuScenes benchmark, where the model is reported under both ST-P3 [stp3] and UniAD [hu2023_uniad] settings. The evaluation metrics include L2 displacement errors at 1, 2, and 3 seconds, along with the average collision rate over the prediction horizon. To further assess the scene understanding ability of OpenDriveVLA, we report its QA prediction performance on three driving visual question answering (VQA) datasets directly after the driving instruction tuning stage, i.e., [nucaption], nuScenesQA [nuscenesqa], and nuX [hintad-nux]. The VQA evaluation results adopt standard NLG metrics, including BLEU, METEOR, CIDEr, BERT-Score, etc.
IV-C Implementation Details
@@
The 3D visual perception module in OpenDriveVLA follows the vision-centric design from [hu2023_uniad], using a ResNet-101 backbone for 2D feature extraction. The perception backbone is pre-trained via multi-task learning on 3D object detection, object tracking, and map segmentation. The resulting BEV feature map has a spatial resolution of $200\times 200$. To construct a unified scene representation, the global SceneSampler applies 2D adaptive pooling to each camera view, subsequently concatenating the pooled multi-view features into a global scene token. Agent and map tokens are extracted from the final layer of their respective QueryTransformer modules. Each token type is then mapped into the language space using a separate two-layer MLP with GeLU activation. We adopt Qwen 2.5-Instruct [qwen2.5] as the pre-trained LLM, which undergoes full parameter tuning during training. Training is performed on 4 NVIDIA H100 GPUs with a batch size of 1, completed in approximately two days. We freeze the 2D backbone during stage 3. During inference, we set the decoding temperature to 0 to ensure deterministic trajectory generation. See supplementary material for detailed training configurations.
IV-D Main Results
Open Loop Trajectory Planning. We evaluate OpenDriveVLA on the open-loop trajectory planning task using both ST-P3 and UniAD metrics, ensuring comprehensive performance assessment across spatial accuracy and collision avoidance. As shown in Table I, OpenDriveVLA achieves state-of-the-art performance across both settings. Specifically, both 3B and 7B version models achieve an average L2 error of 0.33m under ST-P3 metrics, outperforming prior autoregressive language models [mao2024gptdriver, tian2024drivevlm]. On the UniAD metrics, OpenDriveVLA-7B also achieves great performance with an average L2 error of 0.66m. Notably, despite significantly fewer parameters, the 0.5B version still outperforms prior models obviously.
TABLE III: Performance comparison of OpenDriveVLA on the Nu-X dataset [hintad-nux].
Models CIDER BL-4 METEOR ROUGE-L
Hint-UniAD [hintad-nux] 21.7 4.2 12.7 27.0
Hint-VAD [hintad-nux] 22.4 4.2 13.2 27.6
GPT-4o [drivegpt4] 19.0 4.0 10.3 24.9
Gemini 1.5 [geminiteam2024gemini15unlockingmultimodal] 17.6 3.4 9.3 23.4
Vote2CapDETR [vote2cap] 15.3 2.6 10.9 24.2
TOD3Cap [TOD3Cap] 14.5 2.5 10.5 23.5
@@ IV-D Main Results
Driving Question Answering. We access OpenDriveVLA on the driving VQA task across three nuScenes-based datasets (Table II, Table III), reporting results after the second stage of training. OpenDriveVLA reaches best-in-class performance across all three datasets, consistently outperforming previous language-enhanced driving models and general-purpose multimodal baselines among most metrics. On nuCaption dataset, it achieves the best captioning performance among all evaluated models, outperforming both general VLMs LLaVA1.5 [llava1_5] and Mini-GPT4 [zhu2024minigpt4], as well as autonomous driving-specific models such as LiDAR-LLM [nucaption]. For nuScenesQA dataset, OpenDriveVLA also achieves strong performance. Compared to models that directly fuse BEV features with language models such as BEVDet+BUTD [nuscenesqa], it demonstrates clear advantages in object and status-related questions, which highlights the benefit of its spatially grounded visual-language alignment. Notably, the 0.5B version outperforms even the larger 7B on the Nu-X dataset, which shows its powerful scene-understanding ability even with lightweight LLMs.
Figure 3: Visualization of OpenDriveVLA-7B planning actions under original dataset instruction to keep forward (left) and modified instruction to turn right (right). The QA prediction showcases (middle) are from results reported in Table II and Table III. The agent motion prediction results are visualized after the agent-env-ego interaction stage.
IV-E Ablation Study
We conduct ablation studies to evaluate the impact of input modalities and our multi-stage training strategy on OpenDriveVLA’s performance. Additionally, we qualitatively assess the model’s ability to follow diverse driving commands.
Visu Ego Hist Cmd Avg. Collision (%) ↓ Avg. L2 (m) ↓
@@ IV-E Ablation Study
TABLE IV: Ablation study on the effect of different input combinations on OpenDriveVLA-0.5B.
Effect of Input Modalities. We investigate how individual input components contribute to trajectory planning. Table IV presents the results of ablating visual perception, ego state, historical trajectory, and high-level language commands. The inclusion of visual inputs significantly boosts overall performance. Adding textual commands and historical information further improves the predictions, emphasizing the value of semantic intent and temporal context. Notably, ego-state features play a critical role in nuScenes open-loop benchmark, consistent with prior findings [bevplanner].
Effect of Multi-Stage Training Strategy. We evaluate the contribution of each training phase in our staged pipeline incrementally. As shown in Table V, each additional stage consistently improves performance, with the most notable reductions in collision rate observed after Hierarchical Vision-Language Alignment and Agent-Environment-Ego Interaction Modeling. These improvements highlight the effectiveness of cross-modal grounding and interaction-aware reasoning in enhancing safety-critical planning behavior.
Training Stage Avg. Collision (%) ↓ Avg. L2 (m) ↓
1 2 2.5 3 UniAD ST-P3 UniAD ST-P3
0.37 0.13 0.70 0.36
0.32 0.12 0.69 0.35
0.31 0.11 0.68 0.35
0.26 0.09 0.68 0.35
TABLE V: Ablation study on the effect of multi-stage training of 0.5B model. Stage 1, 2, 2.5, and 3 correspond to hierarchical feature alignment, driving instruction tuning, Agent-Env-Ego modeling, and trajectory tuning, respectively.
@@
Effect of Driving Command. Figure 3 presents the qualitative comparison at an intersection under two different driver instructions: keep forward and turn right, with the right turn as the ground truth. OpenDriveVLA accurately adapts its plan to the given command while maintaining context-aware and environment-consistent behavior, demonstrating robust command-following and generalization in complex scenes. In addition, we visualize the QA predictions for the same scene, showcasing the model’s ability to reason over decision-making and traffic scene understanding.
V Conclusion